# Orthogonalisability of joins of graphs

### v1

@@ 1. Introduction
where $q(A)$ denotes the number of distinct eigenvalues of a square matrix $A$, is one of the parameters motivated by Inverse Eigenvalue Problem for Graphs. The study of $q(G)$ was initiated by Leal-Duarte and Johnson in [18], and it has been broadly studied since then (see e.g. [4, 6, 10, 12, 16, 19]).
@@
The minimum number of distinct eigenvalues of joins of graphs has been investigated in several works, e.g., [1, 2, 4, 20]. Recall that the join of two graphs $G$ and $H$, denoted $G\vee H$, is the graph formed by adding all possible edges between $G$ and $H$ to the disjoint union $G\cup H$. It is known that $q(G\vee H)=2$ for all connected graphs $G$ and $H$ such that $\big{|}|G|-|H|\big{|}\leq 2$[21, Theorem 3.4]; see also Ahmadi et al. [4, Theorem 5.2] for the case $G=H$ and Monfared and Shader [23, Theorem 5.2] for $|G|=|H|$. This implies that there does not exist any set of forbidden subgraphs of graphs with $q=2$. It was shown in [21, Example 3.5] that for trees $G$ and $H$ we have $q(G\vee H)=2$ if and only if $\big{|}|G|-|H|\big{|}\leq 2$. In particular, $q(P_{m}\vee P_{n})\geq 3$ for $|m-n|\geq 3$, where $P_{n}$ is the path on $n$ vertices. Recently, joins of disconnected graphs were investigated in [2, 3, 20, 21], where particular attention was given to joins of unions of complete graphs and paths.
@@ 1.1. Notation and terminology
We write =$\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$. For $k\in\mathbb{N}$, we write $[k]=\{1,2,\dots,k\}$. Vectors are written in boldface, and the $i$th component of a vector $\mathbf{v}$ is written as $v_{i}$. For integer tuples $\mathbf{x}\in\mathbb{N}_{0}^{t}$ we write $\iota(\mathbf{x})=|\{i\in[t]:x_{i}=1\}|$ and $|\mathbf{x}|=\sum_{i\in[t]}x_{i}$. If $x_{i}\geq x_{i+1}$ for all $i<t$, then we say that $\mathbf{x}$ is non-increasing. When necessary, we extend integer tuples by appending zero parts at the end, and adopt the useful shorthand $(k^{s}):=(k,k,\dots,k)$, where $k$ appears $s$ times. At times will use this notation for subtuples, too, so that, for example $(4,3^{2},1^{5})=(4,3,3,1,1,1,1,1)$. We use standard notation for conjugate partitions: for $\mathbf{x}\in\mathbb{N}_{0}^{\ell}$, we set $x_{j}^{*}=|\{i\in[\ell]:x_{i}\geq j\}|$ and $\mathbf{x}^{*}=(x_{j}^{*})$. We write $\preceq_{w}$ for the weak majorisation order on non-increasing integer tuples, defined by $\mathbf{x}\preceq_{w}\mathbf{y}$ if and only if $\sum_{i\in[k]}x_{i}\leq\sum_{i\in[k]}y_{i}$ for all $k\in\mathbb{N}$. If in addition $|\mathbf{x}|=|\mathbf{y}|$, then we say that $\mathbf{y}$ majorises $\mathbf{x}$ and write $\mathbf{x}\preceq\mathbf{y}$.
@@
The graphs considered in this work are all finite, simple, undirected graphs with at least one vertex. We say that a graph $G$ is sized by $\mathbf{m}\in\mathbb{N}^{k}$ if $G$ has $k$ connected components with sizes $m_{1},\dots,m_{k}$, respectively, and a pair of graphs $(G,H)$ is sized by a pair $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ if $G$ is sized by $\mathbf{m}$ and $H$ is sized by $\mathbf{n}$. Note that $\iota(\mathbf{n})$ is then the number of isolated vertices of $H$.
@@ 1.2. Overview of the main results
In this work, we investigate pairs $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ for which any pair of graphs $(G,H)$ sized by $(\mathbf{m},\mathbf{n})$ satisfies $q(G\vee H)=2$. Our main results are summarized in the following theorem.
@@ 2. Compatible multiplicity matrices
Let $G$ be a graph with $k$ connected components $G_{1},\dots,G_{k}$. A matrix $V$ with non-negative integer entries fits $G$ if $V$ has $k$ columns and the $i$th column of $V$ has sum $|G_{i}|$ for $1\leq i\leq k$. We say that $V$ is a multiplicity matrix for $G$ if there is a matrix $A\in S(G)$ and real numbers $\lambda_{1}<\dots<\lambda_{r}$, where $r$ is the number of rows of $V$, such that each matrix entry $v_{i,j}$ is the multiplicity of $\lambda_{i}$ as an eigenvalue of $A_{j}$, the submatrix of $A$ corresponding to the connected component $G_{j}$ of $G$. In particular, this implies that the $j$th column of $G$ is an ordered multiplicity list of some matrix in $S(G_{j})$. (The zero entries in the matrix correspond to multiplicity zero, and are ignored in the corresponding ordered multiplicity list of $A_{j}$.) By a 00-$1$ matrix, we mean a matrix whose entries are all in $\{0,1\}$. Monfared and Shader, [22, Corollary 4.3] and [23, Theorem 4.3], showed that any 00-$1$ matrix which fits a graph $G$ is a multiplicity matrix for $G$.
@@
For $r,s\in\mathbb{N}$ and ${\bf r}\in\mathbb{N}_{0}^{r}$, $\mathbf{s}\in\mathbb{N}_{0}^{s}$, we write $\mathcal{A}({\bf r},\mathbf{s})$ for the set of $r\times s$ matrices $X\in\{0,1\}^{r\times s}$ such that $\mathbf{r}$ and $\mathbf{s}$ are the row-sum and the column-sum of $X$, respectively. That is, $X\mathbf{1}_{s}=\mathbf{r}$ and $X^{\top}\mathbf{1}_{r}=\mathbf{s}$.
@@
Let $1\leq k\leq\ell$ and $s\geq 1$. If $(\mathbf{m},\mathbf{n})\in\widetilde{\mathcal{D}}(k,\ell)$ then for any $\mathbf{p}\in\mathbb{N}_{0}^{k}$, $\mathbf{q}\in\mathbb{N}_{0}^{\ell}$ with $|{\bf p}|=|{\bf q}|$, we have $(\mathbf{m}+\mathbf{p},\mathbf{n}+\mathbf{q})\in\widetilde{\mathcal{D}}(k,\ell)$.
@@
Since ${\bf c}\leq(s^{\ell})$ and $|{\bf c}|=ks=|(s^{k})|$, the decreasing rearrangement of ${\bf c}$ is majorised by $(s^{k})$, which has conjugate partition $(s^{k})^{*}=(k^{s})$. By the Gale-Ryser theorem [15, 24] (see also [11, Theorem 2.1.3]), there is an $s\times\ell$ matrix $W\in\mathcal{A}((k^{s}),{\bf c})$. Let $V$ be the $s\times k$ matrix with every entry equal to $1$. Then $V\in\mathcal{A}((k^{s}),(s^{k}))$ and every entry of $V^{\top}W$ is an entry of ${\bf c}\in\mathbb{N}^{\ell}$, so $((s^{k}),{\bf c})\in\widetilde{\mathcal{D}}(k,\ell)$. ∎
@@
Without loss of generality suppose that the entries of $\mathbf{m}$ are arranged in decreasing order. We must show that $m_{k}\geq\psi$, or equivalently, $km_{k}\geq\ell$. There exist ${\bf r}$ and matrices $V\in{\mathcal{A}}({\bf r},\mathbf{m})$, $W\in{\mathcal{A}}({\bf r},\mathbf{n})$ such that $V^{\top}W$ is nowhere zero. By permuting the rows of $V$ and $W$ if necessary, we may assume that the $k$th column of $V$ is $(1^{m_{k}},0,0,\dots)^{\top}$. Since $V$ and $W$ are compatible, each of the $\ell$ columns of $W$ must contain at least one $1$ in the first $m_{k}$ rows; for otherwise, $V^{\top}W$ would have a zero entry. On the other hand, each row-sum $r_{i}$ of $W$ is equal to the corresponding row-sum of $V$, which is at most $k$. Therefore
@@ 3. Join-orthogonalisability, strong and weak suitability
In this section we establish one of our main results, Theorem 3.6, which relates two simple combinatorial conditions to two properties related to orthogonalisability of joins of graphs. We start by defining and discussing these orthogonalisability properties.
@@
By [21, Theorem 3.1 and Corollary 3.3], a pair $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ is always join-orthogonalisable if and only if
@@
Recall [9, Theorem 3.1] that any realisable eigenvalue list of a connected graph $G$ of order $n$ is also realisable for the complete graph $K_{n}$ of the same order. It follows from [21, Theorem 3.1 and Corollary 3.2] that $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ is sometimes join-orthogonalisable if and only if
@@
In light of this, we can state [20, Theorem 4.10] as the following characterisation of sometimes join-orthogonalisability.
@@
holds, then we say that the pair $(\mathbf{m},\mathbf{n})$ is strongly suitable.
We say that a pair $(\boldsymbol{\delta},\boldsymbol{\varepsilon})$ illustrates the weak or strong suitability of $(\mathbf{m},\mathbf{n})$, as appropriate, when the corresponding conditions above hold.
@@
Finally, suppose that $(\mathbf{m},\mathbf{n})$ is weakly suitable, and let $(\boldsymbol{\delta},\boldsymbol{\varepsilon})$ illustrate this. To show that $(\mathbf{m},\mathbf{n})$ is sometimes join-orthogonalisable, it suffices to check that (1) implies that one of the conditions (a), (b) and (c) in Theorem 3.4 holds. First note that $\mathbf{m}-\boldsymbol{\delta}\geq(\psi^{k})$ implies that $|\mathbf{m}-\boldsymbol{\delta}|=k\psi\geq\ell$. Hence we have $|\mathbf{m}|\geq|\mathbf{m}-\boldsymbol{\delta}|\geq\ell$, and $|\mathbf{n}|\geq|\mathbf{n}-\boldsymbol{\varepsilon}|=|\mathbf{m}-\boldsymbol{\delta}|\geq\ell$. So if $\iota(\mathbf{m})=0$, then either (a) or (c) holds. If $\iota(\mathbf{m})\neq 0$, then since $\mathbf{m}\geq\mathbf{m}-\boldsymbol{\delta}\geq(\psi^{k})$, we must have $\psi=1$, so $\ell=k$. Hence $|\mathbf{m}|+\iota(\mathbf{m})=\sum_{i\in[k]:m_{i}\geq 2}m_{i}+2\iota(\mathbf{m})\geq 2k=k+\ell$, so (b) holds. ∎
Determining whether a given pair is weakly or strongly suitable using the definitions of these properties can be inconvenient as it involves as search over various vectors $\boldsymbol{\delta}$ and $\boldsymbol{\varepsilon}$. Our next goal (Proposition 3.9) is to find characterisations of these properties which do not require such a search.
@@
  3. (3)
    clearly follows by (1) and (2).∎
The following lemma is elementary and we omit its easy proof.
@@
    A pair $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ is strongly suitable if and only if it weakly suitable and
@@
Finally, we prove that (3) is equivalent to (4). Without loss, we may assume that $\mathbf{n}$ is non-increasing. Since $n_{1}^{*}=\ell$, observe that $|\mathbf{n}|-n_{2}^{*}-n_{3}^{*}\leq|\mathbf{m}|$ in (3) is equivalent to
@@
If $\mathbf{n}$ satisfies (9), let ${\bf p}=(n_{4}^{*},n_{5}^{*},\dots)^{*}$. Then $|{\bf p}|\leq|\mathbf{m}|-\ell$, and we have $\mathbf{n}=(3^{a},2^{b},1^{c})+{\bf p}$ where $a=n_{3}^{*}$, $b=n_{2}^{*}-n_{3}^{*}$ and $c=\ell-n_{2}^{*}$. In particular, since $\mathbf{n}\leq(3^{\ell})+{\bf p}$, it follows that $\mathbf{n}\preceq_{w}(3^{\ell})+(|\mathbf{m}|-\ell,0^{\ell-1})=(3+|\mathbf{m}|-\ell,3^{\ell-1})$ and so (3) implies (4). Conversely, let $\mathbf{n}\preceq_{w}\mathbf{u}:=(3+|\mathbf{m}|-\ell,3^{\ell-1})$. If $n_{4}^{*}=0$, then (9) clearly holds. Otherwise, let $j=n_{4}^{*}\geq 1$, so that $n_{j}\geq 4$ and $n_{j+1}\leq 3$. Then we have
@@
which implies (9). Thus (4) implies (3). ∎

### v2

@@ 1. Introduction
where $q(A)$ denotes the number of distinct eigenvalues of a square matrix $A$, is one of the parameters motivated by the Inverse Eigenvalue Problem for Graphs. The origins of the study of this parameter can be found in several key texts [10, 13, 18], and many papers have been published on the subject [4, 6, 11, 14, 19, 21, 22].
@@
The minimum number of distinct eigenvalues of joins of graphs has been investigated in several works, e.g., [1, 2, 4, 23]. Recall that the join of two graphs $G$ and $H$, denoted $G\vee H$, is the graph formed by adding all possible edges between $G$ and $H$ to the disjoint union $G\cup H$. It is known that $q(G\vee H)=2$ for all connected graphs $G$ and $H$ such that $\big{|}|G|-|H|\big{|}\leq 2$[24, Theorem 3.4]; see also Ahmadi et al. [4, Theorem 5.2] for the case $G=H$ and Monfared and Shader [26, Theorem 5.2] for $|G|=|H|$. This implies that there does not exist any set of forbidden subgraphs of graphs with $q=2$. It was shown in [24, Example 3.5] that for trees $G$ and $H$ we have $q(G\vee H)=2$ if and only if $\big{|}|G|-|H|\big{|}\leq 2$. In particular, $q(P_{m}\vee P_{n})\geq 3$ for $|m-n|\geq 3$, where $P_{n}$ is the path on $n$ vertices. Recently, joins of disconnected graphs were investigated in [2, 3, 23, 24], where particular attention was given to joins of unions of complete graphs and paths.
@@ 1.1. Notation and terminology
We write =$\mathbb{N}=\{1,2,3,\dots\}$ and $\mathbb{N}_{0}=\{0\}\cup\mathbb{N}$. For $k\in\mathbb{N}$, we write $[k]=\{1,2,\dots,k\}$. Vectors are written in boldface, and the $i$th component of a vector $\mathbf{v}$ is written as $v_{i}$. For integer tuples $\mathbf{x}\in\mathbb{N}_{0}^{t}$ we write $\iota(\mathbf{x})=|\{i\in[t]:x_{i}=1\}|$ and $|\mathbf{x}|=\sum_{i\in[t]}x_{i}$. If $x_{i}\geq x_{i+1}$ for all $i<t$, then we say that $\mathbf{x}$ is non-increasing. When necessary, we extend integer tuples by appending zero parts at the end, and adopt the useful shorthand $(k^{s}):=(k,k,\dots,k)$, where $k$ appears $s$ times. At times we will use this notation for subtuples, too, so that, for example $(4,3^{2},1^{5})=(4,3,3,1,1,1,1,1)$. We use standard notation for conjugate partitions: for $\mathbf{x}\in\mathbb{N}_{0}^{\ell}$, we set $x_{j}^{*}=|\{i\in[\ell]:x_{i}\geq j\}|$ and $\mathbf{x}^{*}=(x_{j}^{*})$. We write $\preceq_{w}$ for the weak majorisation order on non-increasing integer tuples, defined by $\mathbf{x}\preceq_{w}\mathbf{y}$ if and only if $\sum_{i\in[k]}x_{i}\leq\sum_{i\in[k]}y_{i}$ for all $k\in\mathbb{N}$. If in addition $|\mathbf{x}|=|\mathbf{y}|$, then we say that $\mathbf{y}$ majorises $\mathbf{x}$ and write $\mathbf{x}\preceq\mathbf{y}$.
@@
The graphs considered in this work are all finite, simple, undirected graphs with at least one vertex. We say that a graph $G$ is sized by $\mathbf{m}\in\mathbb{N}^{k}$ if $G$ has $k$ connected components with sizes $m_{1},\dots,m_{k}$, respectively, and a pair of graphs $(G,H)$ is sized by a pair $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ if $G$ is sized by $\mathbf{m}$ and $H$ is sized by $\mathbf{n}$. Note that $\iota(\mathbf{n})$ is the number of isolated vertices of $H$.
@@ 1.2. Overview of the main results
In this work, we investigate pairs $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ for which any pair of graphs $(G,H)$ sized by $(\mathbf{m},\mathbf{n})$ satisfies $q(G\vee H)=2$. Our main results are summarised in the following theorem.
@@
a $2P_{2}\vee(P_{2}\cup 2P_{1})$
b $(P_{2}\cup P_{1})\vee(P_{2}\cup 2P_{1})$
Figure 1. The graph in (a) is sized by $(\mathbf{m},\mathbf{n})=((2^{2}),(2,1^{2}))$ which satisfies condition (i) in Theorem 1.1, hence it has $q=2$. The graph in (b) is sized by $(\mathbf{m},\mathbf{n})=((2,1),(2,1^{2}))$ which fails condition (iv) in Theorem 1.1, so this graph has $q>2$.
@@ 2. Compatible multiplicity matrices
Let $G$ be a graph with $k$ connected components $G_{1},\dots,G_{k}$. A matrix $V$ with non-negative integer entries fits $G$ if $V$ has $k$ columns and the $i$th column of $V$ has sum $|G_{i}|$ for $1\leq i\leq k$. We say that $V$ is a multiplicity matrix for $G$ if there is a matrix $A\in S(G)$ and real numbers $\lambda_{1}<\dots<\lambda_{r}$, where $r$ is the number of rows of $V$, such that each matrix entry $v_{i,j}$ is the multiplicity of $\lambda_{i}$ as an eigenvalue of $A_{j}$, the submatrix of $A$ corresponding to the connected component $G_{j}$ of $G$. In particular, this implies that the $j$th column of $G$ is an ordered multiplicity list of some matrix in $S(G_{j})$. (The zero entries in the matrix correspond to multiplicity zero, and are ignored in the corresponding ordered multiplicity list of $A_{j}$.) By a 00-$1$ matrix, we mean a matrix whose entries are all in $\{0,1\}$. Monfared and Shader, [25, Corollary 4.3] and [26, Theorem 4.3], showed that any 00-$1$ matrix which fits a graph $G$ is a multiplicity matrix for $G$.
@@
For $r,s\in\mathbb{N}$ and ${\bf r}\in\mathbb{N}_{0}^{r}$, $\mathbf{s}\in\mathbb{N}_{0}^{s}$, we use the standard notation $\mathcal{A}({\bf r},\mathbf{s})$ for the set of $r\times s$ matrices $X\in\{0,1\}^{r\times s}$ such that $\mathbf{r}$ and $\mathbf{s}$ are the row-sum and the column-sum of $X$, respectively. That is, $X\mathbf{1}_{s}=\mathbf{r}$ and $X^{\top}\mathbf{1}_{r}=\mathbf{s}$.
@@
Let $1\leq k\leq\ell$ and $s\geq 1$. If $(\mathbf{m},\mathbf{n})\in\widetilde{\mathcal{D}}(k,\ell)$, then for any $\mathbf{p}\in\mathbb{N}_{0}^{k}$, $\mathbf{q}\in\mathbb{N}_{0}^{\ell}$ with $|{\bf p}|=|{\bf q}|$, we have $(\mathbf{m}+\mathbf{p},\mathbf{n}+\mathbf{q})\in\widetilde{\mathcal{D}}(k,\ell)$.
@@
Since ${\bf c}\leq(s^{\ell})$ and $|{\bf c}|=ks=|(s^{k})|$, the decreasing rearrangement of ${\bf c}$ is majorised by $(s^{k})$, which has conjugate partition $(s^{k})^{*}=(k^{s})$. By the Gale-Ryser theorem [17, 27] (see also [12, Theorem 2.1.3]), there is an $s\times\ell$ matrix $W\in\mathcal{A}((k^{s}),{\bf c})$. Let $V$ be the $s\times k$ matrix with every entry equal to $1$. Then $V\in\mathcal{A}((k^{s}),(s^{k}))$ and every entry of $V^{\top}W$ is an entry of ${\bf c}\in\mathbb{N}^{\ell}$, so $((s^{k}),{\bf c})\in\widetilde{\mathcal{D}}(k,\ell)$. ∎
@@
Without loss of generality, suppose that the entries of $\mathbf{m}$ are arranged in decreasing order. We must show that $m_{k}\geq\psi$, or equivalently, $km_{k}\geq\ell$. There exist ${\bf r}$ and matrices $V\in{\mathcal{A}}({\bf r},\mathbf{m})$, $W\in{\mathcal{A}}({\bf r},\mathbf{n})$ such that $V^{\top}W$ is nowhere zero. By permuting the rows of $V$ and $W$ if necessary, we may assume that the $k$th column of $V$ is $(1^{m_{k}},0,0,\dots)^{\top}$. Since $V$ and $W$ are compatible, each of the $\ell$ columns of $W$ must contain at least one $1$ in the first $m_{k}$ rows; for otherwise, $V^{\top}W$ would have a zero entry. On the other hand, each row-sum $r_{i}$ of $W$ is equal to the corresponding row-sum of $V$, which is at most $k$. Therefore
@@ 3. Join-orthogonalisability, strong and weak suitability
In this section, we establish one of our main results, Theorem 3.6, which relates two simple combinatorial conditions to two properties related to the orthogonalisability of joins of graphs. We start by defining and discussing these orthogonalisability properties.
@@
By [24, Theorem 3.1 and Corollary 3.3], a pair $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ is always join-orthogonalisable if and only if
@@
Recall [9, Theorem 3.1] that any realisable eigenvalue list of a connected graph $G$ of order $n$ is also realisable for the complete graph $K_{n}$ of the same order. It follows from [24, Theorem 3.1 and Corollary 3.2] that $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ is sometimes join-orthogonalisable if and only if
@@
In light of this, we can state [23, Theorem 4.10] as the following characterisation of sometimes join-orthogonalisability.
@@
holds, then we say that the pair $(\mathbf{m},\mathbf{n})$ is strongly suitable. We say that a pair $(\boldsymbol{\delta},\boldsymbol{\varepsilon})$ illustrates the weak or strong suitability of $(\mathbf{m},\mathbf{n})$, as appropriate, when the corresponding conditions above hold.
@@
Finally, suppose that $(\mathbf{m},\mathbf{n})$ is weakly suitable, and let $(\boldsymbol{\delta},\boldsymbol{\varepsilon})$ illustrate this. To show that $(\mathbf{m},\mathbf{n})$ is sometimes join-orthogonalisable, it suffices to check that (1) implies that one of the conditions (i), (ii) and (iii) in Theorem 3.4 holds. First note that $\mathbf{m}-\boldsymbol{\delta}\geq(\psi^{k})$ implies that $|\mathbf{m}-\boldsymbol{\delta}|=k\psi\geq\ell$. Hence we have $|\mathbf{m}|\geq|\mathbf{m}-\boldsymbol{\delta}|\geq\ell$, and $|\mathbf{n}|\geq|\mathbf{n}-\boldsymbol{\varepsilon}|=|\mathbf{m}-\boldsymbol{\delta}|\geq\ell$. So if $\iota(\mathbf{m})=0$, then either (i) or (iii) holds. If $\iota(\mathbf{m})\neq 0$, then since $\mathbf{m}\geq\mathbf{m}-\boldsymbol{\delta}\geq(\psi^{k})$, we must have $\psi=1$, so $\ell=k$. Hence $|\mathbf{m}|+\iota(\mathbf{m})=\sum_{i\in[k]:m_{i}\geq 2}m_{i}+2\iota(\mathbf{m})\geq 2k=k+\ell$, so (ii) holds. ∎
Determining whether a given pair is weakly or strongly suitable using the definitions of these properties can be inconvenient, as it involves a search over various vectors $\boldsymbol{\delta}$ and $\boldsymbol{\varepsilon}$. Our next goal (Proposition 3.9) is to find characterisations of these properties which do not require such a search.
@@
  3. (iii)
    clearly follows by (i) and (ii).∎
The following lemma is elementary, and we omit its easy proof.
@@
    A pair $(\mathbf{m},\mathbf{n})\in\mathbb{N}^{k}\times\mathbb{N}^{\ell}$ is strongly suitable if and only if it is weakly suitable and
@@
Finally, we prove that (3) is equivalent to (4). Without loss, we may assume that $\mathbf{n}$ is non-increasing. Since $n_{1}^{*}=\ell$, observe that $|\mathbf{n}|-n_{2}^{*}-n_{3}^{*}\leq|\mathbf{m}|$ in (iii) is equivalent to
@@
If $\mathbf{n}$ satisfies (9), let ${\bf p}=(n_{4}^{*},n_{5}^{*},\dots)^{*}$. Then $|{\bf p}|\leq|\mathbf{m}|-\ell$, and we have $\mathbf{n}=(3^{a},2^{b},1^{c})+{\bf p}$ where $a=n_{3}^{*}$, $b=n_{2}^{*}-n_{3}^{*}$ and $c=\ell-n_{2}^{*}$. In particular, since $\mathbf{n}\leq(3^{\ell})+{\bf p}$, it follows that $\mathbf{n}\preceq_{w}(3^{\ell})+(|\mathbf{m}|-\ell,0^{\ell-1})=(3+|\mathbf{m}|-\ell,3^{\ell-1})$ and so (iii) implies (iv). Conversely, let $\mathbf{n}\preceq_{w}\mathbf{u}:=(3+|\mathbf{m}|-\ell,3^{\ell-1})$. If $n_{4}^{*}=0$, then (9) clearly holds. Otherwise, let $j=n_{4}^{*}\geq 1$, so that $n_{j}\geq 4$ and $n_{j+1}\leq 3$. Then we have
@@
which implies (9). Thus (iv) implies (iii). ∎